OpenStudy (superteenblue):

Deidra plotted 3 points from the equation y = 3x + 1 on this coordinate grid. She drew a straight line through the points. Which ordered pair would also be on this line? A. (+5, +11) B. (+6, +18) C. (+3, +10) D. (+5, +13)

OpenStudy (anonymous):

If we plug in x=3 into y=3x+1 we get y=3(3)+1 = 9+1 =10 So,choice C works

OpenStudy (superteenblue):

thank vbro